# ComputerRL: Scaling End-to-End Online Reinforcement Learning for Computer Use Agents

### v1

@@ 5 Future Direction
5.1 Towards Robust Performance
AutoGLM-OS has demonstrated remarkable proficiency across a spectrum of desktop tasks. However, genuine universality requires transcending current boundaries in coverage and generalization. Real-world digital environments are characterized by continual flux and heterogeneity, encompassing unfamiliar applications, emergent workflows, and rare edge cases that lie beyond the scope of existing datasets. A next-generation agent must dynamically adapt to shifting GUIs, unpredictable pop-ups, and entirely novel interfaces. To this end, we are re-architecting data pipelines to facilitate exponential expansion in training diversity and pioneering infrastructure to continuously distill knowledge from real-world user interactions at scale. Furthermore, integrating advanced multimodal perception will empower agents with a deep understanding of both visual and contextual information, enabling them to achieve seamless, human-level adaptability.
5.2 Breakthroughs in Long-Horizon Autonomy
@@
5.3 Foundations for Safe and Aligned Autonomy
@@
6 Conclusion
In this work, we present ComputerRL, a novel computer use agent framework that advances desktop automation through the integration of API-based and GUI actions, efficient data collection, and scalable RL. Our experiments on the OSWorld benchmark demonstrate superior reasoning, accuracy, and generalization compared to prior approaches. This work lays the groundwork for more capable and persistent autonomous computer use agents, highlighting promising directions for future research in intelligent human-computer interaction.
Appendix A Demonstration
@@
A.1 Create Monthly Sales Summary Table
@@
As Figure 6,7 shows, a summary table is created successfully.
A.2 Convert Uppercase Text to Lowercase in Document
@@
A.3 Generate System Resource Monitoring Report
@@
A.4 Copy Image from .xcf File to LibreOffice Writer and Save
@@
A.5 Question Misunderstanding Error
@@
A.6 Click Operation Error

### v2

@@ Appendix F Future Direction
F.1 Towards Robust Performance
AutoGLM-OS has demonstrated remarkable proficiency across a spectrum of desktop tasks. However, genuine universality requires transcending current boundaries in coverage and generalization. Real-world digital environments are characterized by continual flux and heterogeneity, encompassing unfamiliar applications, emergent workflows, and rare edge cases that lie beyond the scope of existing datasets. A next-generation agent must dynamically adapt to shifting GUIs, unpredictable pop-ups, and entirely novel interfaces. To this end, we are re-architecting data pipelines to facilitate exponential expansion in training diversity and pioneering infrastructure to distill knowledge from real-world user interactions at scale continuously.
F.2 Breakthroughs in Long-Horizon Autonomy
@@
F.3 Foundations for Safe and Aligned Autonomy
@@
Appendix G Demonstration
@@
G.1 Create Monthly Sales Summary Table
@@
As Figure 9,10 shows, a summary table is created successfully.
G.2 Convert Uppercase Text to Lowercase in Document
@@
G.3 Generate System Resource Monitoring Report
@@
G.4 Copy Image from .xcf File to LibreOffice Writer and Save
@@
G.5 Question Misunderstanding Error
@@
G.6 Click Operation Error